# Token Reduction Should Go Beyond Efficiency in Generative Models – From Vision, Language to Multimodality

### v2

@@ 4.1 Obtain Informative Visual Representation
MLLMs often suffer from noisy visual inputs that impede fine-grained understanding. We outline key challenges in MLLM visual reasoning: ① Text-Visual Attention Shift: Due to the rotary positional embeddings in LLM decoders, later text tokens disproportionately attend to spatially lower image regions hong2024token , shifting attention away from semantically important areas (e.g., objects at the top of an image); ② Visual Redundancy: Empirical studies lin2025boosting ; wu2024accelerating show that beyond the first few layers, many image tokens contribute little new information, ③ Task-Guided Focus in VQA: In multimodal question answering, the question itself pinpoints relevant image regions (e.g., "kitten color" directs focus to the kitten patch), implying that many image tokens are unnecessary for correct answers song2024less .
Therefore, we position token reduction as a representation-learning optimization: selecting the subset of tokens that preserves informative visual representation. For example, VisPruner zhang2025textvisualattentionexploitingvisual identifies high-value tokens using visual-encoder attention and removes duplicates via clustering to ensure diversity. VTW lin2025boosting observes that visual information migrates into text tokens within early layers; it therefore withdraws all visual tokens after a chosen layer based on KL-divergence criteria. TRIM song2024less leverages the CLIP metric and IQR scoring function to adaptively select image tokens that are crucial for answering questions, while an aggregated token is used to retain additional image information.
@@ 4.2 Better Multimodal Token Alignment
Despite their impressive capabilities, MLLMs continue to face challenges in semantic alignment. Standard vision tokenizers typically split images into fixed-size patches, which can fragment coherent visual entities (e.g., objects or regions) across multiple tokens. This fragmentation weakens the alignment between visual and linguistic representations. Token reduction offers a promising solution by selecting visual tokens based on semantic importance, thereby producing a compact set of tokens that better align with language representations. Specifically, SeTok wu2024towards dynamically clusters visual features into semantically meaningful tokens using a density-peak algorithm, which determines both the number and structure of token groupings per image. This approach preserves both high- and low-frequency semantics, substantially improving concept-level alignment and downstream task performance. M3 cai2024matryoshka introduces a hierarchical token structure that captures coarse-to-fine semantic granularity, allowing different levels of abstraction to be selectively retained depending on task needs.
@@ 4.3 Reduce Overthinking in Reasoning
LLM reasoning. In the context of language models, overthinking refers to generating excessively long or convoluted chains of reasoning that go beyond what is necessary to reach a correct answer. An LLM may produce verbose, repetitive, or even self-contradictory explanations when it fails to converge on a solution-often due to uncertainty sui2025stop ; wang2025harnessing . Such extended reasoning trajectories are inefficient and recent studies show that state-of-the-art reasoners can consume over 15,000 tokens to solve math problems that could be addressed with a concise chain-of-thought (CoT) of just a few hundred tokens hou2025thinkprune . Mitigating overthinking is thus crucial for improving both accuracy and efficiency in generative modeling. By reducing unnecessary tokens during reasoning, LLMs can focus on salient steps, aligning generation with a more concise and logical trajectory.
CoT-Influx huang2023fewer introduces a CoT pruning strategy in which concise reasoning examples are included in the prompt. By pruning unimportant tokens from these examples, more reasoning demonstrations can fit into the context window, surprisingly leading to improved math reasoning accuracy. TokenSkip xia2025tokenskip enables LLMs to skip less important tokens within CoT sequences and learn shortcuts between critical reasoning steps. This allows for controllable CoT compression with adjustable compression ratios, enabling models to automatically trim redundant tokens during reasoning.
MLLM reasoning. MLLMs, which reason over text and other modalities, face similar overthinking issues. In vision-language tasks, overthinking often manifests as excessive processing of visual tokens or overly detailed image descriptions, resulting in inefficiency and potential confusion chen2025zipr1reinforcingtokensparsity . Token reduction techniques in MLLMs aim to promote more focused and sparse reasoning over multimodal inputs. For example, FAST xiao2025fast rewards shorter-than-average token sequences for correct answers, while allowing longer reasoning for more complex tasks. It also adjusts policy optimization constraints to tighten output exploration for simple tasks (thus reducing unnecessary tokens) and loosen it for harder ones to allow deeper reasoning.
@@
4.4 Improve Training Stability
While token reduction has traditionally been employed as a post-training optimization to enhance inference efficiency, recent research indicates its potential to significantly improve training stability when integrated into the pre-training phase gao2023masked ; li2024pruning ; lin2024not , suggesting that selective token utilization during training can lead to more robust and effective model learning.
One notable approach is Rho-1 lin2024not , which involves scoring tokens based on their alignment with a desired distribution using a reference model and then focusing the training loss on tokens with higher scores. Therefore, it effectively filters out noisy or less informative tokens, leading to faster convergence and improved performance. UPFT ji2025first emphasizes the importance of initial reasoning steps in training. By reducing the number of training tokens, UPFT encourages the model to focus on the initial prefix substrings of reasoning trajectories, which are often more stable and contain crucial information. This focus helps the model avoid being influenced by subsequent complex or potentially erroneous information, thereby improving training stability. However, how to design token reduction algorithms that are tightly integrated with training procedures such as in modern methods like GRPO shao2024deepseekmath , remains underexplored. Future research should investigate specialized approaches that incorporate token reduction directly into training objectives, enabling models to learn to prioritize or discard tokens in a task-aware and gradient-aligned manner.
@@ 4.5 Enhance Long Context & Video Understanding
Long-context LLMs. Long-context language modeling presents unique challenges: ① Long texts often contain raw tokens that exhibit repetitive descriptions and irrelevant details that strain the attention mechanism; ② LLM-based agent systems use input data as sequential prompts for reasoning or for switching between multiple tasks, which can lead to overload when the prompt grows too large; ③ It is very difficult to scale up to even longer content for learning more information. Token reduction techniques directly address these issues by distilling extensive input sequences into compact summary vectors or representative tokens. By doing so, models preserve core information such as key events, central themes, or task-specific facts, while significantly decreasing cognitive load. For example, AutoCompressors chevalier2023adapting trains pre-trained LLMs to compress long contexts into compact summary tokens, reducing token length by orders of magnitude to extend context windows and speed up inference. TokenSwift wu2025hours reduces the effective number of tokens that the model dynamically processes during generation by using multi-token parallel generation and n-gram retrieval for token reutilization, therefore enabling efficient ultra-long sequence generation (up to 100K tokens).
Video-based MLLMs. The necessity of token reduction primarily lies in enhancing the model’s effective understanding of video content through: ① Instruction-guided information filtering: token reduction prioritizes selecting visual information relevant to user instructions over raw data volume. ② preserving spatiotemporal structure: token reduction strategically compresses massive spatiotemporal information to retain spatiotemporal dependencies, ensuring the model can capture dynamic semantics, as well as prevent redundant tokens interfere with long temporal reasoning. ③ Preserving semantic integrity: it facilitates feasible processing of extremely long sequences in learning while preserving semantic integrity. ④ Multi-modal alignment: token reduction distills visual information into a compact, semantically aligned form, thereby efficiently bridging the gap between language and vision liu2025hybrid . By doing so, it effectively addresses the challenges posed by the low abstractness and lack of guidance inherent in raw visual inputs, which are the root causes of semantic misalignment and optimization ambiguity in multi-modal models. Recent works illustrate these principles: HICom liu2025hybrid conducts conditional token compression at local and global levels using user instructions as guidance to retain instruction-relevant visual information while reducing computational burden. Video-XL-Pro liu2025video employs reconstructive token compression with a dynamic token synthesizer and semantic-guided masking to generate compact yet comprehensive video tokens for improved MLLM performance and efficiency.
@@ 5 Future Directions
In this section, we propose eight promising directions for token reduction beyond the efficiency benefits, organized into three categories: (i) Algorithmic Innovations (Sec. 5.1$\sim$5.4), (ii) Application Innovations (Sec. 5.5$\sim$ 5.7), and (iii) Hardware-Algorithm Co-Design (Sec. 5.8).
@@ 5.1 Design of New Algorithms
Better Token Importance Metrics. It is critical to re-evaluate how token importance is defined and measured. More robust and unbiased scoring mechanisms can be developed, such as predictors akhauri2025tokenbutler or meta-learning frameworks that go beyond attention-based proxies. These models should capture downstream utility with minimal supervision, enabling adaptive pruning across tasks and domains.
Constructive Token Compression. Token reduction can shift from purely eliminative pruning to strategies that merge spatially or semantically similar tokens into compact summary vectors li2024tokenpacker .
Mitigating Position Bias. In MLLMs, attention-based pruning methods (e.g., FastV) often rely on attention scores from a fixed query token, leading to retained tokens concentrating in specific image regions (e.g., lower corner) wen2025token with potential position bias. Future methods should preserve spatial diversity by enforcing structural uniformity in retained tokens to improve robustness on visual tasks.
Cross-Modal Guided Pruning. Pruning decisions in MLLMs should be guided by inter-modality dependencies, rather than made independently for each modality. For example, text-guided pruning of visual tokens can improve alignment between modalities cao2024madtp . The design should account for joint representations and semantic correspondence across all relevant inputs.
End-to-End Sparsification. Token reduction should consider both the prefill stage and decoding phase for LLMs. This includes dynamically managing the sparsity of KV caches and selectively updating generated tokens, sustaining efficiency gains throughout the entire inference process  huang2024dynamic .
@@
5.2 Token Reduction for Reinforcement Learning
Reinforcement learning (RL)-driven token reduction technology demonstrates significant potential in multimodal large models, with its core lying in balancing computational efficiency and reasoning accuracy through dynamic reward mechanisms and sparsity constraints wang2025beyond . Under the “Fast-Slow Thinking” framework xiao2025fast , RL can guide hierarchical screening of high-value tokens: The fast path employs sparsity strategies (e.g., rule-based rewards or information density scoring) to filter redundant visual or semantic features, while the slow path focuses on refined reasoning. This hierarchical mechanism not only reduces computational overhead by up to 30-50%, but also continuously optimizes token selection strategies via RL’s online feedback.
The sparsity reinforcement paradigm proposed in ZipR1 chen2025zipr1reinforcingtokensparsity highlights RL’s unique advantage in cross-modal generalization. By designing sparsity rewards (e.g., token quantity limits or step compression incentives), models can autonomously generate concise intermediate reasoning chains and even distill such capabilities into lightweight models. For example, in visual question answering tasks, models can dynamically extract key image region features while enhancing answer reliability through multi-candidate contrast mechanisms (e.g., parallel samplings), achieving the goal of “low computational cost, high task accuracy.”
Recent findings further reveal that optimizing only a small subset of high-entropy "forking tokens" can match or even surpass full-gradient updates in reasoning tasks wang2025beyond , suggesting that token-level entropy patterns can guide efficient policy updates. This entropy-aware strategy reinforces the broader view that sparse, selective token refinement is sufficient for preserving reasoning performance in large models. Looking ahead, the integration of such approaches could enhance cross-modal alignment and inference efficiency in real-time and resource-constrained scenarios, supporting a new generation of lightweight yet capable multimodal language agents.
@@ 5.3 From Prompt Tuning to Chain of Thought Reasoning
Current token reduction efforts for prompts have primarily aimed at compressing prompts for efficiency, often with impressive results jiang2023llmlingua ; mu2023learning . Looking forward, token reduction should evolve into enhancing reasoning and maximizing utility per token in context. Instead of focusing solely on making prompts shorter, future research should explore how each remaining token can carry more information or trigger more complex inference during in-context learning and chain of thought reasoning. One direction is to alter the generation paradigm itself, for example, training language models to predict multiple tokens per step gloeckle2024better . Another idea is to enable deeper internal reasoning without increasing prompt length adams2023sparse .
As mentioned in Sec. 4, long CoT chains can become verbose: excessive reasoning steps may introduce errors or obscure logical clarity, particularly in LLM agents where internal reasoning alternates with external tool use gao2025txagent ; wang2024toolgen . Token reduction may serve a critical role in this context by compressing intermediate reasoning into a compact representation. For example, approaches based on next-token prediction mu2023learning ; zhang2025lightthinker can distill intermediate thinking chains into a set of dense, information-rich tokens. These compressed representations can then replace the full intermediate context and serve as inputs for subsequent reasoning steps.
@@ 5.4 Complementary to Other Methods
Token reduction can complement other efficiency techniques, such as quantization. By selectively reducing the number of tokens processed during inference, models can improve both performance and efficiency, particularly when paired with quantization strategies li2025mergevq . Traditional key-value cache quantization methods often suffer from accuracy loss due to their inability to handle outlier tokens that carry distinct or rare features. To mitigate this issue, Outlier Token Tracking su2025accuratekvcachequantization identifies outlier tokens during decoding and excludes them from quantization, preserving full-precision necessary representations and improving key-value cache quantization accuracy. Similarly, Agile-Quant shen2024agile incorporates token pruning as a preprocessing step to reduce the impact of activation outliers. It prunes tokens based on their attention to the start-of-sequence token, discarding those with low attentiveness, which often appear in adjacent channels and contribute to quantization noise. This targeted pruning reduces interaction distances between salient tokens and helps maintain model accuracy under low-bit quantization settings.

### v3

@@ 4.1 Obtain Informative Visual Representation
MLLMs often suffer from noisy visual inputs that impede fine-grained understanding. We outline key challenges in MLLM visual reasoning: ① Text-Visual Attention Shift: Due to the rotary positional embeddings in LLM decoders, later text tokens disproportionately attend to spatially lower image regions Hong et al. (2024), shifting attention away from semantically important areas (e.g., objects at the top of an image); ② Visual Redundancy: Empirical studies Lin et al. (2025); Wu et al. (2024) show that beyond the first few layers, many image tokens contribute little new information, ③ Task-Guided Focus in VQA: In multimodal question answering, the question itself pinpoints relevant image regions (e.g., "kitten color" directs focus to the kitten patch), implying that many image tokens are unnecessary for correct answers Song et al. (2025).
Therefore, token reduction can serve as a representation-learning optimization: selecting the subset of tokens that preserves informative visual representation. For example, VisPruner Zhang et al. (2025c) identifies high-value tokens using visual-encoder attention and removes duplicates via clustering to ensure diversity. VTW Lin et al. (2025) observes that visual information migrates into text tokens within early layers; it therefore withdraws all visual tokens after a chosen layer based on KL-divergence criteria. TRIM Song et al. (2025) leverages the CLIP metric and IQR scoring function to adaptively select image tokens that are crucial for answering questions, while an aggregated token is used to retain additional image information.
@@ 4.2 Better Multimodal Token Alignment
Despite their impressive capabilities, MLLMs continue to face challenges in semantic alignment. Standard vision tokenizers typically split images into fixed-size patches, which can fragment coherent visual entities (e.g., objects or regions) across multiple tokens. This fragmentation weakens the alignment between visual and linguistic representations. Token reduction offers a promising solution by selecting visual tokens based on semantic importance, thereby producing a compact set of tokens that better align with language representations. Specifically, SeTok Wu et al. (2025b) dynamically clusters visual features into semantically meaningful tokens using a density-peak algorithm, which determines both the number and structure of token groupings per image. This approach preserves both high- and low-frequency semantics, substantially improving concept-level alignment and downstream task performance. M3 Cai et al. (2025) introduces a hierarchical token structure that captures coarse-to-fine semantic granularity, allowing different levels of abstraction to be selectively retained depending on task needs.
@@ 4.3 Reduce Overthinking in Reasoning
LLM reasoning. In the context of language models, overthinking refers to generating excessively long or convoluted chains of reasoning that go beyond what is necessary to reach a correct answer. An LLM may produce verbose, repetitive, or even self-contradictory explanations when it fails to converge on a solution-often due to uncertainty Sui et al. (2025); Wang et al. (2025a). Such extended reasoning trajectories are inefficient and recent studies show that state-of-the-art reasoners can consume over 15,000 tokens to solve math problems that could be addressed with a concise chain-of-thought (CoT) of just a few hundred tokens Hou et al. (2025). This issue is particularly acute in LLM agents, where internal reasoning alternates with external tool use Gao et al. (2025a); Wang et al. (2024b); excessive steps can obscure logical clarity and lead to error accumulation. Mitigating overthinking is thus crucial. By trimming unnecessary tokens, LLMs can focus on salient steps, aligning generation with a more concise trajectory.
CoT-Influx Huang et al. (2024b) introduces a CoT pruning strategy in which concise reasoning examples are included in the prompt. By pruning unimportant tokens from these examples, more reasoning demonstrations can fit into the context window, surprisingly leading to improved math reasoning accuracy. TokenSkip Xia et al. (2025) enables LLMs to skip less important tokens within CoT sequences and learn shortcuts between critical reasoning steps. This allows for controllable CoT compression with adjustable compression ratios, enabling models to automatically trim redundant tokens during reasoning.
MLLM reasoning. MLLMs, which reason over text and other modalities, face similar overthinking issues. In vision-language tasks, overthinking often manifests as excessive processing of visual tokens or overly detailed image descriptions, resulting in inefficiency and potential confusion Chen et al. (2025). Token reduction techniques in MLLMs aim to promote more focused and sparse reasoning over multimodal inputs. For example, FAST Xiao et al. (2025a) rewards shorter-than-average token sequences for correct answers, while allowing longer reasoning for more complex tasks. It also adjusts policy optimization constraints to tighten output exploration for simple tasks (thus reducing unnecessary tokens) and loosen it for harder ones to allow deeper reasoning.
@@
4.4 Improve Training Stability & Efficiency
While token reduction has traditionally been employed as a post-training optimization to enhance inference efficiency, recent research indicates its potential to significantly improve training stability when integrated into the pre-training phase Gao et al. (2023); Li et al. (2025c); Lin et al. (2024), suggesting that selective token utilization during training can lead to more robust model learning.
One notable approach is Rho-1 Lin et al. (2024), which involves scoring tokens based on their alignment with a desired distribution using a reference model and then focusing the training loss on tokens with higher scores. Therefore, it effectively filters out noisy or less informative tokens, leading to faster convergence and improved performance. UPFT Ji et al. (2025) emphasizes the importance of initial reasoning steps in training. By reducing the number of training tokens, UPFT encourages the model to focus on the initial prefix substrings of reasoning trajectories, which are often more stable and contain crucial information. This focus helps the model avoid being influenced by subsequent complex or potentially erroneous information, thereby improving training stability.
Additionally, integrating token reduction with training procedures like GRPO Shao et al. (2024) is gaining traction; for instance, recent work reveals that optimizing only a subset of high-entropy ”forking tokens” matches full-gradient updates Wang et al. (2025b), suggesting that entropy patterns can effectively guide efficient policy learning. Future research should investigate specialized approaches that incorporate token reduction directly into training objectives, enabling models to learn to prioritize or discard tokens in a task-aware and gradient-aligned manner.
@@ 4.5 Enhance Long Context & Video Understanding
Long-context LLMs. Long-context language modeling presents unique challenges: ① Long texts often contain raw tokens that exhibit repetitive descriptions and irrelevant details that strain the attention mechanism; ② LLM-based agent systems use input data as sequential prompts for reasoning or for switching between multiple tasks, which can lead to overload when the prompt grows too large; ③ It is very difficult to scale up to even longer content for learning more information. Token reduction techniques directly address these issues by distilling extensive input sequences into compact summary vectors or representative tokens. By doing so, models preserve core information such as key events, central themes, or task-specific facts, while significantly decreasing cognitive load. For example, AutoCompressors Chevalier et al. (2023) trains pre-trained LLMs to compress long contexts into compact summary tokens, reducing token length by orders of magnitude to extend context windows and speed up inference. TokenSwift Wu et al. (2025d) reduces the effective number of tokens that the model dynamically processes during generation by using multi-token parallel generation and n-gram retrieval for token reutilization, therefore enabling efficient ultra-long sequence generation (up to 100K tokens).
Video-based MLLMs. The necessity of token reduction primarily lies in enhancing the model’s effective understanding of video content through: ① Instruction-guided information filtering: token reduction prioritizes selecting visual information relevant to user instructions over raw data volume. ② preserving spatiotemporal structure: token reduction strategically compresses massive spatiotemporal information to retain spatiotemporal dependencies, ensuring the model can capture dynamic semantics, as well as prevent redundant tokens interfere with long temporal reasoning. ③ Preserving semantic integrity: it facilitates feasible processing of extremely long sequences in learning while preserving semantic integrity. ④ Multi-modal alignment: token reduction distills visual information into a compact, semantically aligned form, thereby efficiently bridging the gap between language and vision Liu et al. (2025e). By doing so, it effectively addresses the challenges posed by the low abstractness and lack of guidance inherent in raw visual inputs, which are the root causes of semantic misalignment and optimization ambiguity in multi-modal models. Recent works illustrate these principles: HICom Liu et al. (2025e) conducts conditional token compression at local and global levels using user instructions as guidance to retain instruction-relevant visual information while reducing computational burden. Video-XL-Pro Liu et al. (2025a) employs reconstructive token compression with a dynamic token synthesizer and semantic-guided masking to generate compact yet comprehensive video tokens for improved MLLM performance and efficiency.
@@ 5 Future Directions
[…]
@@ 5.1 Design of New Algorithms
Better Token Importance Metrics. It is critical to re-evaluate how token importance is defined and measured. More robust and unbiased scoring mechanisms can be developed, such as predictors Akhauri et al. (2025) or meta-learning frameworks that go beyond attention-based proxies. These models should capture downstream utility with minimal supervision, enabling adaptive pruning across tasks and domains.
Constructive Token Compression. Token reduction can shift from purely eliminative pruning to strategies that merge spatially or semantically similar tokens into compact summary vectors Li et al. (2024a).
Mitigating Position Bias. In MLLMs, attention-based pruning methods (e.g., FastV) often rely on attention scores from a fixed query token, leading to retained tokens concentrating in specific image regions (e.g., lower corner) Wen et al. (2025) with potential position bias. Future methods should preserve spatial diversity by enforcing structural uniformity in retained tokens to improve robustness on visual tasks.
Cross-Modal Guided Pruning. Pruning decisions in MLLMs should be guided by inter-modality dependencies, rather than made independently for each modality. For example, text-guided pruning of visual tokens can improve alignment between modalities Cao et al. (2024). The design should account for joint representations and semantic correspondence across all relevant inputs.
End-to-End Sparsification. Token reduction should consider both the prefill stage and decoding phase for LLMs. This includes dynamically managing the sparsity of KV caches and selectively updating generated tokens, sustaining efficiency gains throughout the entire inference process  Huang et al. (2025).
@@ 5.2 From Prompt Tuning to Chain of Thought Reasoning
Current token reduction efforts for prompts have primarily aimed at compressing prompts for efficiency, often with impressive results Jiang et al. (2023); Mu et al. (2023). Looking forward, token reduction should evolve into enhancing reasoning and maximizing utility per token in context. Instead of focusing solely on making prompts shorter, future research should explore how each remaining token can carry more information or trigger more complex inference during in-context learning and chain of thought reasoning. One direction is to alter the generation paradigm itself, for example, training language models to predict multiple tokens per step Gloeckle et al. (2024). Another idea is to enable deeper internal reasoning without increasing prompt length Adams et al. (2023).
As mentioned in Sec. 4, long CoT chains can become verbose: excessive reasoning steps may introduce errors or obscure logical clarity, particularly in LLM agents where internal reasoning alternates with external tool use Gao et al. (2025a); Wang et al. (2024b). Token reduction may serve a critical role in this context by compressing intermediate reasoning into a compact representation. For example, approaches based on next-token prediction Mu et al. (2023); Zhang et al. (2025b) can distill intermediate thinking chains into a set of dense, information-rich tokens. These compressed representations can then replace the full intermediate context and serve as inputs for subsequent reasoning steps.
@@
5.3 Efficient Reasoning with Reinforcement Learning
Reinforcement-learning (RL)-driven token reduction has shown strong promise for improving reasoning efficiency in both LLMs and MLLMs. The key challenge is to balance compute and reasoning quality via dynamic reward design, sparsity-inducing constraints, and adaptive control of effective token length Aggarwal and Welleck (2025); Luo et al. (2025); Wu et al. (2025c).
In language reasoning, length awareness is incorporated either by diversifying reasoning formats during pre-training Su et al. (2025a) or by adding explicit length penalties in the RL stage Arora and Zanette (2025); Hou et al. (2025). While effective, most approaches implicitly assume static task complexity or rely on hand-specified length constraints, which can be suboptimal under heterogeneous workloads. A natural next step is adaptive reasoning: using RL to learn per-instance budget allocation from intrinsic task difficulty NVIDIA (2025); Qwen (2025). In parallel, performing reasoning in a compressed latent space can yield substantial computational savings Shen et al. (2025b), but current methods often degrade due to poorly structured latent representations. A promising direction is to inject explicit logical structures into the latent space to enable more controllable and compositional reasoning Helff et al. (2025).
Beyond language modality, under the “Fast-Slow Thinking” framework Xiao et al. (2025a), RL can supervise hierarchical selection of high-value tokens: a fast branch applies sparsity signals (e.g., rule-based rewards or information-density scoring) to prune redundant visual/semantic features, while a slow branch allocates computation to refined reasoning. Additionally, RL enables a Think-with-Image paradigm Su et al. (2025d) by letting models adapt visual granularity: VisionThink Yang et al. (2025b) adopts a progressive resolution strategy, using RL to selectively request high-resolution inputs only for fine-grained tasks like OCR. PixelThink Wang et al. (2025c) addresses visual overthinking in segmentation by adjusting reasoning length based on task difficulty. These methods demonstrate how RL can dynamically calibrate both input resolution and reasoning depth across various tasks. Looking ahead, the integration of such approaches could enhance cross-modal alignment and inference efficiency in real-time and resource-constrained scenarios, supporting a new generation of lightweight yet capable multimodal language agents.
@@ 5.4 Complementary to Other Methods
Token reduction can complement other efficiency techniques, such as quantization. By selectively reducing the number of tokens processed during inference, models can improve both performance and efficiency, particularly when paired with quantization strategies Li et al. (2025b). Traditional key-value cache quantization methods often suffer from accuracy loss due to their inability to handle outlier tokens that carry distinct or rare features. To mitigate this issue, Outlier Token Tracking Su et al. (2025c) identifies outlier tokens during decoding and excludes them from quantization, preserving full-precision necessary representations and improving key-value cache quantization accuracy. Similarly, Agile-Quant Shen et al. (2024) incorporates token pruning as a preprocessing step to reduce the impact of activation outliers. It prunes tokens based on their attention to the start-of-sequence token, discarding those with low attentiveness, which often appear in adjacent channels and contribute to quantization noise. This targeted pruning reduces interaction distances between salient tokens and helps maintain model accuracy under low-bit quantization settings.